wat is the answer for 5 5/6 + 2 2/3
Answer by stanbon(75887)

5 5/6 + 2 2/3
-------
= (35/6) + (8/3)
----
= (35/6) + (16/6)
----
= 51/6
-----
= 8 1/2
==================
Cheers,
Stan H.